Did I solve this problem correctly for the thickness of aluminum foil that is 50cm long and 20cm wide, and 5.4 grams?
H=LW/V
H=50*20/5.4
H=1000/5.4
H=185cm.

Answer by chessace(471)

That would be one thick foil!
The key is to use units!
Info is missing!
Vol = HLW cm^3
Density = ?? gram(g) per cm^3; Not given (or is it 5.4 g/cm^3 and the weight is not given?).
weight = vol * density(D) = HLW * D in cm^3 *(g/cm^3) = g
H = Weight/(LWD); g / (cm cm (g/cm^3)) = cm
H = 5.4 / (50 * 20 * D[??] )
H = 0.0054 / D[??] cm